Errata for Mathematical Writing, second printing (last updated 29 April 2019) (the first printing contained several errors that were corrected in the second)

[On the copyright page about the MAA Notes series, delete the bottom line that says “A version of these notes… ”.] Page 4, the last line of rule 22 when it should be when the proper word is Page 5, replacement for lines −3 and −2

∃_{n_{0}∈N_{0}} ∋ ∀_{p∈P} p ≥ n_{0} ⇒ f(p) = 0

then there exists a bijection N_{1} → L^{+}(P, N_{0}) such that if n ↦ f then

Page 8, line 18 k(b_{i} − b_{j}) k(b_{j} − b_{i}) Page 9, line 9 alternate spellings alternative spellings Page 9, line 13 some of solutions some of the solutions Page 10, line 14 should omitted should be omitted Page 11, line 21 sentence; No sentence. No Page 12, line 11 from the bottom (c_{i} − c_{j} + 1) remans as a unit c_{i} − c_{j} + 1 remains as a unit Page 14, line 19 at the authors request at the authors’ request Page 15, line 12 close them”. close them.” Page 15, line 5 from the bottom that had been that it had been Page 15, line 4 from the bottom Concrete Math draft Concrete Maths draft Page 17, line 7 from the bottom (r/2 − k)’. (r/2 − k)’.) Page 20, line 5 from the bottom letters in variable letters in variable-Page 21, line 3 fixed width type fixed-width type Page 24, line 11 idea idea idea Page 26, line 3 piece of code piece of software Page 26, line 14 from the bottom good and gave a nice intuition of good, and said that it gave a nice intuition about Page 26, line 7 from the bottom phases can replicate places can replicate Page 26, line 2 from the bottom sistences sistencies Page 27, line 2 from the bottom had called a certain variable called ‘scan’ had given a certain variable the name ‘scan’ Page 29, line 1

enter ‘Type of message (1-6):’ enter data by the words ‘Type of message (1-6):’.

Page 31, line 8 now famous now-famous Page 31, line 11 journal”, journal,” Page 31, line 17 accepted” replied accepted,” replied Page 31, line 4 from the bottom the 1960’s 1960s Page 33, line 6 from the bottom (and most damning indictment) (the most damning indictment) Page 34, line 12 some few journals a few journals Page 34, line 13 can strongly effect can strongly affect Page 35, line 9 refereees referees Page 37, line 8 It an extremely It is an extremely Page 37, line 16 Algorithmic translation Algebraic translation Page 38, line 13 Grimm’s Grimms’ Page 38, line 16 Grimm’s Grimms’ Page 38, line 4 from the bottom line-width line width Page 39, line 1 Quantatative Quantitative Page 39, line 7 from the bottom asynchonous asynchronous Page 42, line 9 from the bottom From 1900–1960 From 1900 to 1960 Page 46, bottom line must be spec(α) must be in spec(α) Page 47, line 11 malheur a‘ malheur à Page 47, line 11 instruire! instruire ! Page 47, line 15 agir: Il est l’aime agir : il est l’âme Page 48, line 9 from the bottom Grimm’s Grimms’ Page 49, line 15 from the bottom a little about the the book a little about the book Page 49, line 14 from the bottom Concrete Math Concrete Mathematics Page 49, line 8 from the bottom and Holidays and holidays Page 50, line 23 back from Mr. Flanagan back from Mr. Flanagan Page 51, line 6 from the bottom hoard horde Page 52, line 24 Academie Académie Page 53, line 9 Fine, said, Don Fine, said Don Page 54, line 6 punchline punch line Page 55, line 2 (June 1987) is (June 1987), is Page 55, lines 13 and following [These lines needn’t be indented.] Page 59, line 11 research Journal research journal Page 59, line 16 from the bottom by by by Page 60, line 2 bleachers.’ ” bleachers.’ ”) Page 60, line 5 of Analysis in Analysis Page 60, line 11 Introduction to The Theory An Introduction to the Theory Page 60, line 15 Mathematical Principles of Mathematical Page 64, line 7 to get thee to get the Page 68, line 4 to starts to start Page 71, line 4 Ginsburg Ginsberg Page 72, at the bottom [See Lamport’s article “How to write a proof,” American Mathematical Monthly 102 (1995), 600–608.] Page 73, line 17 from the bottom word-processor wordprocessor Page 73, line 16 from the bottom pin-pointing pinpointing Page 73, line 9 from the bottom organised—and got an A– organised; and he got an A^{−} Page 75, line 7 from the bottom asked what was for asked what it was for Page 77, line 2 from the bottom Greek boys: Taking Greek boys: taking Page 78, line 15 from the bottom to the word count and the word count Page 78, line 10 from the bottom better not be the your better not be your Page 78, line 7 from the bottom the the the Page 79, line 18 from the bottom glue: Words glue: words Page 80, line 3 aware how aware of how Page 80, line 5 fountain pen.” fountain pen.”) Page 84, line 12 relator realtor Page 90, line 5 from the bottom Seirpiński Sierpiński Page 90, line 2 from the bottom write up write-up Page 91, line 15 from the bottom readership”. readership.” Page 92, line 20 Chigaco Manual Chicago Manual Page 100, line 22 they were too). they were, too.) Page 101, line 4 Grimm’s Grimms’ Page 102, line 1 Grimm’s Grimms’ Page 105, line 2 vertices”. vertices.” Page 105, line 17 quality”. quality.” Page 108, line 10 from the bottom balderdash”. balderdash.” Page 109, line 21 so fas as so far as Page 110, line 7 “the reader and I” ‘the reader and I’ Page 111, line 17 this thing?”) was this thing?”) was Page 111, bottom line said said Paul. Page 114, line 5 the the people the people Page 114, line 8 ‘höffentlich’ ‘hoffentlich’ Page 114, line 4 from the bottom leaving out a that leaving out a ‘that’ Page 114, bottom line repeat it twice repeat it again Page 115, line 21 before a which before a ‘which’ Page 116, new material for the end Postscript about “God is in the details” (see page 48): William Safire’s column on Language in the New York Times and the International Herald Tribune, July 31, 1989, discusses this mysterious phrase as well as its counterpart, “The Devil is in the details.” Nobody has been able to trace either one to a definite source. Safire cites Shapiro who claims to cite Nietzsche—but without chapter and verse. Safire also says that Justin Kaplan, editor of Bartlett’s Familiar Quotations, is searching too. According to Kaplan, “Flaubert has been suggested, but nobody can find it in his writings.”

Perhaps the following facts will be helpful. A biography of Mies by Franz Schulze (Chicago, 1985) has a relevant footnote on page 281:

The aphorism, “God is in the details,” has been endlessly attributed to Mies, though I have found no one who ever heard him say it. In Meaning in the Visual Arts (New York, 1955, p. 5), Erwin Panofsky quotes Flaubert: “Le bon Dieu est dans le détail.”

Actually Schulze should have referred not to page ‘5’, but to page ‘v’—where Panofsky drops Flaubert’s name but gives no hint of location. David Spaeth, another Mies biographer, answered a query from Don as follows on December 6, 1985:

The statement “God is in the details” was made by Mies. It [only] appeared in print in an article by Peter Blake entitled “The difficult art of simplicity” which was published in Architecture Forum vol. 108, May 1958, pages 126–131. However, it was a statement Mies made in class many times. George Danforth, one of Mies’ very first students, remembers Mies saying it in the early 1940s.

Paul Roberts (PMR) wrote to Nigel Rees, who produces a BBC radio show based on quotation identification, asking for his opinion. Rees found the phrase mentioned in the New York Herald Tribune obituary of Mies, 1969, and said that Mies “certainly popularized it even if he didn’t originate it.” PMR also turned up a significant clue that might in fact be the true origin of the saying. At least it carries things back further than anyone else has been able to do so far: On pages 13, 14, 229, and 286 of E. H. Gombrich’s 1970 biography of Aby Warburg, a prominent historian of Renaissance art, Gombrich states that Warburg used the phrase “Der liebe Gott steckt im Detail” as the motto of a seminar series at Hamburg University in the fall of 1925; he also says that Warburg often repeated this motto. Thus, Mies probably learned the phrase in Germany. But was it original with Warburg? We can’t rule out Flaubert and Nietzsche until their complete works have been made available in electronic form.

Particulars, as every one knows, make for virtue and happiness; generalities are intellectually necessary evils.

— Aldous Huxley, Brave New World.